OpenStudy (anonymous):

Express y as a function of x. What is the domain? log4y = (x+log4)/4

OpenStudy (anonymous):

Why is the answer y = 10^x and domain is real numbers?

OpenStudy (anonymous):

writing the equation in exponential form you have: \(\large 10^{\frac{x+log4}{4}}=4y \)

OpenStudy (anonymous):

how'd you know it would be 10...

OpenStudy (anonymous):

logs written without a base is log base 10..

OpenStudy (anonymous):

oh yeah

OpenStudy (anonymous):

and i'm assuming the left side of your equation is \(\large log(4y) \) and not \(\large log_4y \)

OpenStudy (anonymous):

yes you're correct

OpenStudy (anonymous):

ohh shoot, it's only: Log4y = x + log4

OpenStudy (anonymous):

no... \(\large log(4y)=log4+logy \) if you're simplifying the left side of your equation first..

OpenStudy (anonymous):

I meant the original question is log4y = x + log4... what happened to the x?

OpenStudy (anonymous):

that last post I put is just simplifying the left side of your original equation... is this you original equation: \(\large log(4y)=\frac{x+log4}{4} \) or \(\large log(4y)=x+log4 \)

OpenStudy (anonymous):

ohh.. the second one

OpenStudy (anonymous):

i'm guessing it is the second one and not the first because you said the answer is y=10^x.... this makes it much simpler.... the original equation is: \(\large log(4y)=x+log4 \) so using the properties of logs on the left side of this equation you have log(4y) = log4 + logy so the equation is now: \(\large log4+logy=x+log4 \) cancelling log4 from both sides you have: \(\large \color {red}{log4}+logy=x+\color {red}{log4} \) so \(\large logy=x \) now this can be written in exponential form: \(\large y=10^x \)

OpenStudy (anonymous):

I don't understand this part: log(4y) = log4 + logy

OpenStudy (anonymous):

there is a property (a very useful one): \[\huge log_b(M\cdot N)=log_bM + log_bN \]

OpenStudy (anonymous):

ohh I know that one but what happened to the x-- you replaced it with log y..

OpenStudy (anonymous):

ohh, never mind. I get it now, lol thank you! :D

OpenStudy (anonymous):

ok.. let's start from the beginning with the original, correct equation: \[\large \color {red}{log(4y)}= x+log4\] using the property on the left side of the equation we get: \[\large \color {red}{log4+logy}= x+log4\] you ok so far?

OpenStudy (anonymous):

yes I'm okay with it so far

OpenStudy (anonymous):

do I transpose log 4 to the right side?

OpenStudy (anonymous):

yes... subtract log4 from both sides of the equation, you have \(\large logy=x \) left. now just write this equation out in exponential form to get \(\large 10^x = y\)

OpenStudy (anonymous):

ohh, by the way.. when do I divide logs? I know that if I subtract logs then I divide them correct?

OpenStudy (anonymous):

the property you're referring to is this: \[\huge log_b(\frac{M}{N})=log_bM-log_bN \]

OpenStudy (anonymous):

the log of a quotient(division) is the difference of logs....

OpenStudy (anonymous):

yes that one. So I only divide when they're both different logs?

OpenStudy (anonymous):

"difference" meaning subtraction....

OpenStudy (anonymous):

yeah I know, so in this case the log4-log4 -- I can divide?

OpenStudy (anonymous):

log4 is just some number.... so a number minus itself is zero...

OpenStudy (anonymous):

ohh.. so when exactly do I use the rule about dividing differences..?

OpenStudy (anonymous):

and yeah you're right too: log4 - log4 = log(4/4) = log(1) = 0

OpenStudy (anonymous):

ohhh, okay. thank you!!

OpenStudy (anonymous):

just a longer way to the simplified answer...

OpenStudy (anonymous):

yw...:)